# GSTAR: Gaussian Surface Tracking and Reconstruction

### v2

@@ 4.3 Ablations
Scene flow warping. For each frame initialization, we construct scene flow to warp the Gaussian Surfaces from the previous frame, effectively handling large deformations. This approach significantly enhances tracking quality and prevents the optimization process from getting trapped in local minima.
@@ 5 Discussion
Limitations. GSTAR may face challenges with complex or sudden topology changes, such as when a new person suddenly enters the scene. Its dependence on multi-view video data restricts its applicability in general public scenarios. Like other Gaussian-based methods, it requires substantial resources for Gaussian optimization and mesh optimization, though it benefits from real-time rendering.
Conclusions. We propose GSTAR, a unified method for high-quality appearance reconstruction, surface reconstruction, and 3D tracking. Our approach represents dynamic surfaces by binding Gaussians to mesh faces. For surfaces with changing topology, new surfaces are reconstructed by unbinding Gaussians. GSTAR effectively handles a wide range of dynamic scenes, paving the way for new applications of Gaussian-based representations.

### v3

@@ 4.3 Ablations
Scene flow warping. For each frame initialization, we construct scene flow to warp the Gaussian Surfaces from the previous frame, effectively handling large deformations. This approach significantly enhances tracking quality and helps prevent the optimization process from getting trapped in local minima.
@@ 5 Discussion
Limitations. GauSTAR may face challenges with complex or sudden topology changes, such as when a new person suddenly enters the scene. Its dependence on multi-view video data restricts its applicability in general public scenarios. Transparent and specular surfaces pose challenges for most surface reconstruction methods, particularly for Gaussian splitting. We leave handling such surfaces for future work.
Conclusions. We propose GauSTAR, a unified method for high-quality appearance reconstruction, surface reconstruction, and 3D tracking. Our approach represents dynamic surfaces by binding Gaussians to mesh faces. For surfaces with changing topology, new surfaces are reconstructed by unbinding Gaussians. GauSTAR effectively handles a wide range of dynamic scenes, paving the way for new applications of Gaussian-based representations.
Acknowledgements
This work was partially supported by the Swiss SERI Consolidation Grant ”AI-PERCEIVE”. The authors thank all the participants of the captured datasets. Part of the computations were performed on the ETH Zürich Euler cluster.